Circular No. 7562

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

XTE J0052–723 R. Corbet, Goddard Space Flight Center (GSFC) and Universities

Space Research Association; F. E. Marshall, GSFC; and C. B. Markwardt, GSFC and University of Maryland, report: “We have discovered a new transient x-ray pulsar in the direction of the Small Magellanic Cloud from RXTE Proportional Counter Array observations made on 2000 Dec. 27 and 2001 Jan. 5. Pulsations were seen with a period of 4.782 ± 0.001 s and a double-peaked pulse profile. Subsequent slew observations on Jan. 9 give a source location of α = 0^{h}52^{m}11^{s}, δ = −72^{o}20′.3 (equinox 2000.0) with a 90-percent statistical uncertainty of ∼ 2′ in α and 1′ in δ. Some additional uncertainty in the source position may be present, due to source variability. The mean source intensity on Jan. 9 was ∼ 8 mCrab (2–10 keV). Earlier observations of this region on 2000 Dec. 13 and 20 did not detect this source.”

SUNGRAZING COMETS Further to IAUC 7560, D. Hammer reports his measurements for ad-

ditional Kreutz sungrazing comets found by J. Danaher (C/2000 A2 in C3 data), M. Oates (Nov. comets and C/2000 X5 in C2 data), X.-m. Zhou (C/2000 X4 in C2 data), T. Lovejoy (C/2000 Y4 in C2 and C3 data), and M. Boschat (remaining Dec. comets in C2 data; both Dec. 5 comets also visible in C3) on SOHO website images. The reduced observations and orbital elements by B. G. Marsden are given on the MPECs cited below.

Comet  2000 UT  α_{2000}  δ_{2000}  MPEC
C/2000 A2 Jan. 14.596  19^{h}57.^{m}6  −23^{o}08′  2001-A32
C/2000 W4 Nov. 19.388  15 32.3  −22 52  2001-A32
C/2000 W5  27.288  16 14.1  −22 55  2001-A35
C/2000 X1 Dec. 3.518  16 43.1  −24 00  2001-A43
C/2000 X2  5.904  16 53.6  −24 53  2001-A43
C/2000 X3  5.654  16 50.8  −26 01  2001-A43
C/2000 X4  5.435  16 51.7  −24 14  2001-A47
C/2000 X5  7.076  16 59.3  −24 29  2001-A47
C/2000 X6  7.451  17 03.4  −24 35  2001-A47
C/2000 X7  10.643  17 16.4  −24 50  2001-A47
C/2000 Y4  17.263  17 46.6  −27 34  2001-A47
2001 January 14  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2001 CBAT  Daniel W. E. Green